OpenStudy (britnee_paige99):

a metal right cylinder is melted and recast as a sphere. The radius of the cylinder was 4 and the height was 18. Find the volume and surface area of the sphere.

OpenStudy (wolf1728):

Volume of cylinder = pi * radius^2 * height Volume of cylinder = pi * 4^2 * 18 Volume of cylinder = 904.779 Volume of a sphere = (4/3) * pi * r^3 We need the sphere volume to equal 904.778 so that means we have to solve for the radius r^3 = Volume of a sphere / (4/3) * pi r^3 = 904.779 / (4/3) * pi r^3 = 216 r = 6 Surface Area of a Sphere = 4 * pi * r^2 Surface Area = 452.3893

OpenStudy (anonymous):